LEARNING TEXT FROM THE WEB Advanced Data Analysis Project

Stella M. Salvatierra

Advisory Committee Stephen Fienberg, Tom Mitchell and Brian Junker

April 16, 1998

EXAMPLE OF THE WEB PAGE DATABASE
faculty:
http://www.cs.cornell.edu/Info/Department/Annual95/Faculty/VanLoan.html
http://www.cs.cornell.edu/Info/People/skeshav/
http://www.cs.cornell.edu/Info/People/rdz/rdz.html
.
.
course:
http://www.cs.cornell.edu/Info/Courses/Current/CS212/CS212.html
http://www.cs.cornell.edu/Info/Courses/Current/CS611/home.html
http://www.cs.cornell.edu/Info/Courses/Spring-96/CS211/CS211.html
.
.
EXAMPLE OF THE CONNECTIVITY DATABASE
(#:g22435 *novalue*
     (underlined.text "Om Damani")
     (column 20)
     (line 55)
     (url "http://www.cs.utexas.edu/users/damani")
        (link.range texas-om-p.-damani)
        (link.domain texas-cs304p---teaching-assistants---spring-1996)
     (generalizations (u.connection)))
(#:g22443 *novalue*
     (underlined.text "1994-5 Annual Report")
     (column 6)
     (line 22)
     (url "http://www.cs.utexas.edu/users/UTCS/report/1995/")
        (link.range texas-contents)
        (link.domain texas-utcs-publications)
     (generalizations (u.connection)))
.
.

100 MOST FREQUENT WORDS
 [1] "edu"               "office"            "rainbowthreedigit"
 [4] "rainbowtwodigit"   "rainbowdigits"     "rainbowonedigit"
 [7] "rainbowtime"       "are"               "in"
[10] "other"             "can"               "also"
[13] "be"                "by"                "students"
[16] "course"            "home"              "page"
[19] "algorithms"        "this"              "the"
[22] "for"               "and"               "we"
[25] "will"              "an"                "of"
[28] "data"              "as"                "design"
[31] "with"              "from"              "some"
[34] "information"       "links"             "more"
[37] "homework"          "you"               "mail"
[40] "cs"                "rainbowphonenum"   "is"
[43] "that"              "to"                "all"
[46] "one"               "not"               "class"
[49] "lecture"           "may"               "at"
[52] "have"              "on"                "software"
[55] "program"           "last"              "rainbowfourdigit"
[58] "university"        "computer"          "science"
[61] "department"        "fall"              "programming"
[64] "group"             "language"          "new"
[67] "your"              "here"              "or"
[70] "if"                "email"             "time"
[73] "postscript"        "student"           "lab"
[76] "project"           "am"                "using"
[79] "it"                "professor"         "system"
[82] "which"             "work"              "performance"
[85] "phone"             "he"                "about"
[88] "engineering"       "systems"           "based"
[91] "computing"         "web"               "has"
[94] "research"          "my"                "me"
[97] "parallel"          "distributed"       "rainbowfivedigit"
[100] "conference"        "interests"

"NON INFORMATIVE" WORDS
[1] "edu"   "are"   "in"    "other" "can"   "also"  "be"    "by"    "this"
[10] "the"   "for"   "and"   "we"    "will"  "an"    "of"    "as"    "with"
[19] "from"  "some"  "more"  "you"   "mail"  "cs"    "is"    "that"  "to"
[28] "all"   "not"   "may"   "at"    "have"  "on"    "your"  "or"    "if"
[37] "email" "am"    "it"    "which" "he"    "about" "web"   "has"

CONFUSION MATRICES BASED ON 57 WORDS

• RAINBOW (Souped up Naive Bayes Classifier)

Rates nearly equivalent to those using 2,000 words Classified as

True Students Projects Faculties Courses Totals Students  144  4  10  5  163 Projects  2  16  7  0  25 Faculties  9  0  32  1  42 Courses  8  8  3  54  73 Totals  163  28  52  60  303

Students Projects Faculties Courses
Accuracy  88%  64%  76%  74%
Coverage  88%  57%  62%  90%

• Linear Discriminant Model (four classes)

Classified as

True Students Projects Faculties Courses Totals Students  159  1  6  1  167 Projects  10  15  0  1  26 Faculties  10  2  34  0  46 Courses  20  1  0  52  73 Totals  199  19  40  54  312

Students Projects Faculties Courses
Accuracy  95%  58%  74%  71%
Coverage  80%  79%  85%  99%

• Mixed Nested Model (Logistic [student vs other] and LDA [remaining three classes]) Classified as

True Students Projects Faculties Courses Totals Students  152  5  9  0  166 Projects  6  9  8  3  26 Faculties  6  2  32  6  46 Courses  2  2  3  66  73 Totals  166  18  52  75  311

Students Projects Faculties Courses
Accuracy  92%  33%  69%  90%
Coverage  91%  52%  6%1  88%